# SingVisio: Visual Analytics of Diffusion Model for Singing Voice Conversion

### v1

@@ 5.5 Metric View
6 Evaluation
This section details the evaluation approach for the SingVisio, aimed at understanding its effectiveness in demonstrating the processes of the diffusion model. The evaluation is carried out through a structured user study, designed to assess both objective understanding and subjective experiences of the users. Given the novel nature of SingVisio, it stands as a unique tool in the field of audio generation, thus a direct comparison with existing systems is not feasible. As such, the focus of the evaluation is on the system’s standalone capabilities in usability, facilitating analysis and affecting users’ perception of the singing voice conversion process.
6.1 User Study Set-up
The user study of SingVisio was designed to gather insights into how effectively the system aids users in visualizing and analyzing the complex process and outcomes of the diffusion model for SVC.
Participants A diverse group of individuals participated in the study. We recruited 20 participants (P1-P20), including students and professionals with backgrounds in machine learning, signal or music processing, and related fields. Their ages ranged from 20 to 35 years, with varied levels of experience in machine learning and music and singing processing, ranging from less than one year to more than three years. This mix of participants ensured a broad perspective on the system’s performance across different user groups.
Questionnaire The questionnaire was designed to capture both objective and subjective aspects of the SingVisio. It comprised 36 questions, among which 17 were objective, as listed in Table 1 in A; 14 were subjective, shown in Fig. 9.
  • 1.
    Objective Questions (O1-O17) evaluate the participants’ understanding of SingVisio’s functions and the insights they can derive about the process of diffusion model from its use. These questions are directly related to the five analytical tasks (T1-T5) previously detailed in Section 4.2, assessing the participants’ grasp of the system’s analytical capabilities.
  • 2.
    Subjective Questions (S1-S14) focus on gathering participants’ overall impressions and evaluations of SingVisio. It is structured around four key aspects: explainability, analysis functionality, design effectiveness, and usability. These aspects were selected based on the recommendations by Rossi  et al. Rossi et al. (2010), ensuring a comprehensive evaluation framework that aligns with established user experience principles. The responses are recorded using a 5-point Likert scale, ranging from 1 (strongly disagree) to 5 (strongly agree).
Procedure The user study was structured into several sessions. Participants were first oriented with a comprehensive tutorial about SingVisio, ensuring they were familiar with its functions and capabilities. They were then asked to engage with the system to try out its key features during the hands-on session. Following this, participants completed the formal assessment in the form of an online questionnaire, which included both objective and subjective queries. Finally, an optional feedback section was provided for any additional comments or suggestions.
6.2 Results and Analysis
The completion time for the questionnaire averaged 1 hour 44 minutes ($\mu=6227.36$, $\sigma=6521.96$). The objective evaluation showed an average accuracy of 85.88% ($\sigma=0.35$), while the subjective evaluation averaged 4.44 out of 5 ($\sigma=0.79$). Fig. 8 and  9 provide a detailed summary of the objective and subjective evaluation results, respectively.
Fig. 8: The results of objective evaluation. The histogram shows the accuracy of each question.
6.2.1 Objective Evaluation
The analysis of users’ objective responses informs the discussion of analytical tasks in Section 4.2. Overall, the average accuracy is high (85.88%). Among the five analytical tasks, T1 recorded the highest accuracy at 91.25% on average, with T3 following at 90.00%. In contrast, T5 gains the lowest accuracy at 78.33% on average.
Step-wise Diffusion Generation Comparison (T1). We designed four questions (O1-O4) related to the diffusion generation process in SVC. These questions involve (1) the transition of Mel spectrograms from noisy to clear, (2) the F0 curve transitions from blurred to clear, and (3) the improvement in the quality of generated audio from unintelligibility to intelligibility, enabling users to understand and perceive the audio generation process in the diffusion model. From the objective results shown in Fig. 8, we can see that the accuracy ranking for these four questions is: O4, O3, O1/O2. Q4 received the highest score ($\mu=100.00\%$, $\sigma=0$), indicating that our system effectively facilitates user understanding of the quality of the converted singing voice. It’s worth noting that the result of Q4 can be obtained through both visual observations of the Mel spectrogram and auditory perception of the audio, highlighting the characteristic of our system that combines both visual and auditory elements. The high accuracy of O3 ($\mu=95.00\%$, $\sigma=0.22$) indicates that our system allows users to clearly observe the F0 range of the audio. The scores for O1/O2 ($\mu=85.00\%$, $\sigma=0.36$) demonstrate the effectiveness of our system for the understanding and analysis of the diffusion generation process at each step.
Step-wise Metric Comparison (T2). For task T2, we designed four questions encompassing: (1) understanding the model evaluation metrics themselves (O6), (2) the relationship between evaluation metrics and model performance (O5), (3) the trend of metric changes during the diffusion conversion process (O7), and (4) the iterative generation paradigm of the diffusion model implied by this trend (O8). The accuracy ranking for these questions is O7, O8, O5, O6. The highest score for O7 ($\mu=95.00\%$, $\sigma=0.22$) indicates that our system obviously demonstrates the relationship between evaluation metrics and model performance, which also forms the foundation for understanding T2, underpinning the validity and reliability of the outcomes for other questions in T2. O8, with the second-highest score, shows that our system effectively allows users to understand the relationship between changes in evaluation metrics and the underlying diffusion model generation paradigm, validating our system design for visualizing the diffusion generation mechanism.
O6, a relatively simpler concept about the model evaluation metrics themselves, scored the lowest. Users reported unfamiliarity with the metric, although we provided a help button in the top right corner of the metric view, displaying a metric tips dialog box with definitions, calculation methods, and relationships to model performance. We analyzed two reasons for O6’s low score: users might not have thoroughly reviewed the system introduction before the user study, which detailed system operation guides; alternatively, the help button might not have been sufficiently prominent, leading to users overlooking or being unaware of it. To address this, we further improved the system by setting the button to a question mark shape, making the help button more prominent and user-friendly.
Pair-wise SVC Comparison with Different Target Singers (T3). Regarding this task, three questions were designed, encompassing: (1) the influence of the target singer on the outcome of the SVC, where differences in timbre among various target singers are the main focus. To aid user understanding, the system provides audio references from different target singers for evaluative listening, with O9 addressing the discernibility of timbre differences; (2) Additionally, since F0 and timbre are correlated, the system presents the corresponding F0 contour to help users better grasp the distinctiveness of target singers, and in O10, we assess whether the system enables users to compare the F0 range differences in the reference audio of different target singers; (3) Finally, O11 deals with the comparison of SVC results under different target singers during the diffusion generation process, touching both on the understanding of the SVC task and the diffusion model.
The accuracy ranking for these questions is O9/O10, O11. O9 and O10 both achieved a 95.00% accuracy rate, indicating the system’s effectiveness in comparing the differences between various target singers; O11 had an 80.00% accuracy rate, which is comparatively lower. We believe this issue is related to the user’s academic and professional background. The question involves the technical term “harmonic”, defined as horizontal bright lines in the Mel spectrogram. For users with a purely musical background (P11), it’s challenging to distinguish the differences in terms of spectrum.
Pair-wise SVC Comparison with Different Source Singers (T4). T4 aims to analyze and understand SVC under different source conditions. For T4, three questions were designed: (1) the timbre differences between two different sources (O12); (2) The differences of the motion trajectories of diffusion steps with two different source melody conditions (O13); (3) The patterns in timbre and fundamental frequency range of the final conversion results (at diffusion step 0) when converting two different sources to the same target singer (O14).
The accuracy ranking for these questions is O12, O14, O13. O12 scored the highest ($\mu=90.00\%$, $\sigma=0.30$), indicating that the system effectively displays the timbre differences between different sources. O14 scored the second-highest ($\mu=85.00\%$, $\sigma=0.36$), showing the system’s effectiveness in demonstrating conversions of two different sources to the same target. O13 scored the lowest ($\mu=75.00\%$, $\sigma=0.43$), with most incorrect answers concentrated on option 2, i.e., “It is unclear what the motion trajectories look like.” Upon revisiting participant P20, who chose this incorrect option, and after further discussion, it was revealed that there was a misunderstanding regarding option 2 of O13. When asked again, they provided the correct answer. Overall, the results reflect that users can understand the differences in SVC under various source conditions through our system, affirming its effectiveness for completing T4.
Pair-wise SVC Comparison with Different Source Songs (T5). In cases where the singing voices of the same singer performing different songs are used as two distinct sources, when converted to the same target, the timbre in the converted results is the same, but the content of the songs is different. T5 aims to analyze this scenario in SVC. To this end, three questions were designed: (1) verify if the system can demonstrate that singing voices of the same singer with different songs have different contents (O15); (2) and (3) assess if the system can aid users in understanding the specific details in terms of the timbre (O16) and content (O17) in the conversion results, when converting two sources with different content from the same singer to the same target.
The accuracy ranking for these questions is O16, O15/O17. O16 has the highest accuracy ($\mu=85.00\%$, $\sigma=0.36$), indicating the system effectively displays the detailed information for the SVC, such as the delivering of timbre in conversion results. However, the accuracy in content (O17) is relatively lower. The most incorrect choices in O17 were for option 1 “The singing content in Target (Singer 13)”. Participant P11, a music student, unfamiliar with the task of the SVC, understandably chose incorrectly. However, P16 and P17, doctoral students in the field of audio generation, were revisited and it was found they had not read the question carefully, overlooking that it asked about two conversion results, not just one, and that there is a corresponding relationship between the two results and the options provided. After understanding the question clearly, they made the correct choice. The accuracy for O15 and O17 is the same ($\mu=75.00\%$, $\sigma=0.43$). We believe the users who chose incorrectly in O15 did not take the task seriously, as a comparative listening of the singing voice audios presented by the system would reveal whether the contents in the audio are consistent or not.
Fig. 9: Description and results of subjective questionnaires. The leftmost column lists four assessment aspects. The rightmost column shows mean ± standard deviation.
6.2.2 Subjective Evaluation
We conduct subjective evaluations of four aspects, including explainability, functionality, effectiveness, and usability. Overall, the assessment results yield an average score of 4.44 ($\sigma=0.79$). Notably, it achieved the highest score in analysis functionality ($\mu=4.65$, $\sigma=0.53$) and also performed well in explainability ($\mu=4.46$, $\sigma=0.77$). Detailed results for each dimension and question are presented in Fig. 9.
Explainablility The primary goal in designing SingVisio is to use visual analysis techniques to demystify the black-box nature of complex neural network-based models and enhance the interpretability of diffusion models, in the context of singing voice conversion tasks. As shown in Fig. 9, the subjective assessment results across four dimensions indicate that explainability scored the second highest, demonstrating the effectiveness of SingVisio in interpreting diffusion models and singing voice conversion.
To validate explainability of SingVisio, we designed four questions: (1) Whether the system facilitates to compare diffusion results at different steps, which is the foundation for the other three questions (S1); (2) Whether the metrics curves aid in understanding changes during the diffusion generation process (S2); (3) Whether comparing converted results under two different source conditions helps in deep understanding the singing voice conversion task (S3); (4) Whether the system is helpful in understanding the working mechanism of the iterative generation process of the diffusion model (S4).
Among these questions, S3 scored the highest ($\mu=4.65$, $\sigma=0.47$), indicating the system’s effectiveness in aiding users to understand and explain SVC under different conditions. S1, S2, and S4 received identical high scores ($\mu=4.65$), further confirming that our system provides a comprehensive understanding and explanation for the diffusion model in the context of the SVC task.
A user specializing in speech enhancement research mentioned that SingVisio is particularly helpful for beginners, “From a beginner’s perspective, SingVisio greatly aids in understanding the diffusion model process. The automatically playing Mel spectrogram changing from noisy to clean with diffusion step evolution, vividly demonstrates the diffusion generation process. Moreover, the system clarifies the relationship between evaluation metrics and model performance, enabling the understanding of how different metrics impact different facets of the synthesized voice. Comparing singing voice conversions across different source singers, target signers, and songs further validates which metrics relate to content and which to timbre or gender. For a beginner, this system has been immensely educational about singing voice conversion.”
Analysis Functionality The test results reveal that in the subjective assessment across four dimensions, the score in the analysis functionality dimension is the highest ($\mu=4.65$, $\sigma=0.53$). This dimension’s subjective evaluation is designed to verify the system’s support for analysis across tasks T1-T5 and includes four specific questions: (1) Whether the system supports interactive selection and comparison of results at different diffusion steps (S5); (2) Whether the system displays metrics curves that change with diffusion steps (S6); (3) Whether the system supports comparison of converting the same source to two different target singers (S7); (4) Whether the system supports visualization of intermediate results in the diffusion model (S8).
Among these questions, S5 and S6 scored the highest ($\mu=4.70$, $\sigma=0.45$), with all users agreeing or strongly agreeing that SingVisio supports the comparison of generated results at different diffusion steps (T1) and the analysis of evaluation metrics (T2). S7, designed for the analysis tasks T3-T5, received agree and strongly agree from all but one neutral user, indicating that SingVisio effectively supports analysis for T3-T5.
Visual Design Effectiveness To validate the effectiveness of our system design, we formulate four questions, encompassing: (1) Whether the Projection View effectively aids in observing the dynamic motion trajectory of diffusion steps and in controlling them (S9); (2) The effectiveness of the Step View in understanding the entire diffusion generation process (S10); (3) Whether the Comparison View effectively displays useful information across five modes (S11); (4) An overarching question: the effectiveness of the interactive design of the SingVisio (S12).
All four questions scored above 4 points, indicating that all users agree or strongly agree that our views’ design and the system’s interactive design are effective. S10 received the highest score ($\mu=4.65$, $\sigma=0.47$), and S11 gained the second highest ($\mu=4.35$, $\sigma=0.85$), demonstrating that Step View and Comparison View, designed specifically for T1-T5, are effective.
Usability Participants in the user study included those with over three years of experience in machine learning and signal processing, some new to these fields, and others with a purely musical background. More than two-thirds of these users strongly believed in the user-friendly user interfance and ease of learning of SingVisio (S13). Over half of the users strongly recommended SingVisio to others who could benefit from its use (S14). This demonstrates that our system is user-friendly for diverse users, regardless of their background. Particularly noteworthy is the feedback from a student specializing in signal processing, who found the visualization of SingVisio in audio generation extremely interesting and innovative. SingVisio helped him to understand audio generation from a new perspective, and he is highly motivated to recommend this system to his research group.
@@ 7 Case Study
E1 showed great interest in understanding the principles and mechanisms of diffusion models, as well as how to use diffusion models for singing voice conversion. Upon opening the SingVisio system, Step View at the top displays an animation showing the Mel spectrogram transitioning from pure noise to clarity, alongside audio that became progressively intelligible. E1 commented that this dynamic display intuitively demonstrated the entire process of singing voice conversion, making the generation process more interpretable and comprehensible. Moreover, E1 mentioned that listening to the voice transition from one blurred timbre to another clear timbre was quite fascinating.
The overall evaluation from E2 was very positive. E2 described the system as a comprehensive and user-friendly visualization tool for diffusion models in SVC tasks that allows for both an overview and a detailed study. E2 praised the system’s interactivity, particularly enjoying the ability to see the changes in the Mel spectrogram at different steps by moving the mouse, clearly showcasing the entire diffusion process. Furthermore, E2 liked the metric comparison feature, stating that I used to know only that higher metrics are better, lower metrics are worse, and what they specifically represent. But this is the first time I’ve directly observed the fluctuations of different metrics throughout the diffusion process, from the starting point to the gradually stabilizing endpoint, and I could also hear the sounds corresponding to different steps.
@@ 8 Expert Study
We invite two domain experts (E3 and E4) to participate in an expert study to evaluate the system based on its usability and effectiveness. They were not involved in the system design process, nor did they participate in the user study and case study. E3 is a researcher who has been engaged in machine learning and voice conversion research for more than 3 years. E4 is also a researcher primarily focusing on singing voice conversion and is strongly interested in XAI.
8.1 Procedure
We carried out an expert study, engaging with each expert individually in semi-structured interviews that spanned 60 minutes. Initially, we presented our research’s goal, the architecture of diffusion-based SVC, and our visualization tool, i.e., SingVisio. Subsequently, we encouraged the experts to examine and assess the system’s features. We then gathered their opinions on the system’s general usability and user interactions. Our findings and the feedback from the experts are summarized below.
8.2 System Usefulness
Both experts acknowledged SingVisio as a valuable tool for validating domain knowledge. They observed that the system clearly demonstrates each step’s results during the data generation process in the diffusion model; E3 said that it vividly displays the denoising process of the diffusion model, showing the Mel spectrogram gradually becoming clearer, and the converted audio progressively moving from unintelligible to intelligible.
E3 also stated, “The system facilitates the comparison of results among different source speakers, target speakers, and songs, generating and displaying corresponding conversion results for straightforward comparison.” E3 also suggested incorporating a tooltip feature in the metric view, which would display descriptions of each metric when the mouse hovers over them, ensuring users clearly understand the metrics. However, there is already a help button in the top right corner of the metric view. Clicking this button opens a dialog box with metric tips, explaining the meanings of the metrics and their relationship to model performance. Upon being informed of this existing feature, E3 agreed that this design choice was satisfactory.
8.3 Visual Designs and Interfactions
E3 and E4 both agree that the system exhibits excellent interactivity. In Control Panel, they could select different source singers, songs, and target singers. Comparison View allowed them to easily compare the differences among various source singers, songs, and target singers. Furthermore, Step View clearly displays the generated results under different conditions. Overall, the system conveniently facilitated their exploration and comparison. E3 stated, “This feature enables users to intuitively assess the differences in conversion performance between different speakers and songs, empowering them to make informed decisions based on their specific needs.”
8.4 Insight and Inspiration
Both domain experts believe the system provides them with valuable insights. E3 mentioned that SingVisio aids in understanding the working principles of the diffusion model for SVC deeply. E4 highlighted that for practitioners in the voice conversion field, especially in singing voice conversion, by exploring comparisons of source songs, source singers, and target singers, we can observe that the diffusion model’s learning capability varies across different samples. For instance, it performs poorly with samples having very low F0. This insight significantly informs us about which samples are challenging for the model and which are easier.

### v2

@@ 5.5 Metric View
5.6 Implementation Details
The web application is designed to provide an interactive and user-friendly interface for visualizing spectrogram differences. It uses D3.js and TailwindCSS for the front-end, ensuring a clean and dynamic user interface. Specifically, D3.js handles the visualization, allowing for detailed and interactive spectrogram comparisons. TailwindCSS ensures a responsive and aesthetic design, enhancing user experience. The back-end is powered by Flask and Gunicorn, enabling efficient dynamic step sampling and efficient data retrieval with multiple workers. Specifically, Flask serves as the core framework, managing API requests and data processing. Gunicorn operates as the WSGI HTTP server, providing concurrency through multiple workers for fast data retrieval and processing. This architecture ensures that the application is both robust and scalable, capable of handling real-time spectrogram analysis and visualization efficiently.
Mel spectrogram and Fundamental Frequency Contour (F0 contour) are extracted from the audio using a signal processing algorithm. Mel spectrogram is a 2D representation with the dimensions of Time*Channel, where the time axis captures the progression of the audio signal over time, and the channel axis represents the frequency components or Mel bins, providing a comprehensive view of the signal’s spectral content. The Mel spectrogram is color-coded to indicate the intensity or magnitude of different frequencies over time. Bright colors, such as yellow and red, represent high energy or the presence of specific frequencies, while darker colors represent lower energy or the absence of those frequencies. F0 contour is a key concept in the fields of speech processing and music analysis, especially in the study of prosody, intonation, and melody. It refers to the variation in the pitch of a voice over time. The fundamental frequency, or F0, is the lowest frequency of a periodic waveform and determines the pitch of the sound, which is one of the primary auditory attributes used to distinguish different sounds in speech and music. This contour line may be drawn as a continuous curve that rises and falls to depict changes in F0. The F0 contour line is colored red in this work, to distinguish it from the Mel spectrogram.
@@ 6 Case Study
E1 focused on the step view, observing the transition of the Mel spectrogram from noisy to clean. Initially, the Mel spectrogram appeared chaotic with many random patterns, lacking clear, distinct features. As the process continued from step 999 to step 0, the spectrogram gradually became clearer, displaying sharp lines representing fundamental frequencies and harmonics. Correspondingly, the initial audio sounded indistinct and lacked clarity, presenting hissing and other unwanted sounds. Eventually, the vocals became well-defined and easy to discern, with almost no unwanted sounds or interference. E1 commented that this dynamic display intuitively demonstrated the entire process of SVC, making the generation process more interpretable and comprehensible. Moreover, E1 mentioned that listening to the voice transition from one blurred timbre to another clear timbre was quite fascinating.
E2 mainly interacted with the system by dragging the step axis to control the diffusion reverse step, observing the differences in the generated results at various steps. E2 also focused on the metric view. E2 clicked on the help button shaped like a question mark in the top right corner of the metric view (as shown in Fig. 2) to learn about the definitions of metrics and their correlation with model performance. E2 then clicked on the MCD metric bar, prompting the system to show five Metric Curves over Diffusion Steps in the Comparison View. E2 moved the mouse over the MCD metric curve and the system displayed the corresponding MCD value, Mel spectrogram and audio of the corresponding step. Additionally, E2 listened to the corresponding audio at different steps, providing an audible perception of the changes. E2 observed that all metric values changed from the starting point to a gradually stabilizing endpoint throughout the diffusion process. E2 mentioned that this was the first time they directly observed the fluctuations of metrics throughout the diffusion process. E2 described the system as a comprehensive and user-friendly visualization tool for diffusion models in SVC tasks that allows for both an overview and a detailed study.
@@ 7 Expert Study
We invite two domain experts (E3 and E4) to participate in an expert study to evaluate the system based on its usability and effectiveness. They were not involved in the system design process, nor did they participate in the user study and case study. E3 is a researcher who has been engaged in machine learning and voice conversion research for more than 3 years. E4 is also a researcher primarily focusing on SVC and is strongly interested in XAI.
System Usefulness Both experts acknowledged SingVisio as a valuable tool for validating domain knowledge. They observed that the system clearly demonstrates each step’s results during the data generation process in the diffusion model. Specifically, in a Mel spectrogram, noise appears as random speckles or fuzzy areas. Early in the reverse diffusion process (step 999), the spectrogram has high noise levels because the model is just starting to refine the audio. By step 50, the noise decreases, resulting in a cleaner spectrogram. This indicates successful noise reduction and improved audio quality. Harmonic structures, seen as horizontal lines and spectral patterns show the distribution of energy across frequencies.
Visual Designs and Interactions From the t-SNE visualization of projection embedding (such as step embedding) in the projection view, experts observe a distinct pattern transitioning from a decentralized to a more centralized structure (as illustrated in Fig 2). In the reverse process of a diffusion model, each step builds on the output of the previous step to remove noise. Viewed as an optimization problem, each step minimizes the difference between the original data and the current estimate. As this process progresses, the latent representations increasingly resemble the original data points, causing them to appear more clustered in the t-SNE plot.
Insight and Inspiration Both domain experts believe the system provides valuable insights. They observed the transformation of the Mel spectrogram from noise to a clear signal during the diffusion generation process in SVC. It was found that when the target singer’s F0 is low and dense, more steps are required for the signal to become clear, indicating greater difficulty in converting to such target singers. Frequent and dense F0 changes increase modeling complexity, necessitating more steps to accurately generate these variations while avoiding distortion and maintaining harmonic structure consistency. Fine-tuning model parameters for low and dense F0 cases can yield better results. Additionally, increasing the quantity and diversity of such data can enhance model robustness and generalization capability.